# exampleframelinecolor=gray!30, linewidth=0.5pt, backgroundcolor=gray!5, roundcorner=2pt, innertopmargin=8pt, innerbottommargin=8pt, innerleftmargin=8pt, innerrightmargin=8pt, skipabove=skipbelow=

### v2

@@ D.4 Proof of Theorem 4: Rank-Stability of ABBA Initialization
For any input $x\in\mathbb{R}^{d_{\text{in}}}$ with i.i.d. entries, the output of the ABBA adapter is given by:
$y_{i}=s_{\text{ABBA}}\sum_{j=1}^{d_{\text{in}}}M_{1,ij}M_{2,ij}\cdot x_{j}\quad\forall i\in\{1,2,\ldots,d_{\text{out}}\}.$ (14)
The second moment can then be expressed as:
@@
We next compute the gradients with respect to the input:
$(\nabla_{x})_{j}$ $=\sum_{i=1}^{d_{\text{out}}}\frac{\partial L}{\partial y_{i}}\frac{\partial y_{i}}{\partial x_{j}}=\sum_{i=1}^{d_{\text{out}}}\frac{\partial y_{i}}{\partial x_{j}}\cdot(\nabla)_{y_{i}}.$
From Equation (14), we have:
$\frac{\partial y_{i}}{\partial x_{j}}=s_{\text{ABBA}}\cdot M_{1,ij}M_{2,ij}.$
Substituting, we obtain:
$(\nabla_{x})_{j}=s_{\text{ABBA}}\sum_{i=1}^{d_{\text{out}}}M_{1,ij}M_{2,ij}\cdot(\nabla)_{y_{i}}.$
Backward $\text{2}^{\text{nd}}$ moment.
We now compute the second moment of the gradient:
$\mathbb{E}[(\nabla_{x})_{j}^{2}]=s_{\text{ABBA}}^{2}\cdot\mathbb{E}\left[\left(\sum_{i=1}^{d_{\text{out}}}M_{1,ij}M_{2,ij}\cdot(\nabla_{y})_{i}\right)^{2}\right].$
Given our initialization, the entries of $M_{1}$ and $M_{2}$ are i.i.d., and thus:
$\mathbb{E}[y_{i}^{2}]$ $=s_{\text{ABBA}}^{2}\cdot\sum_{j=1}^{d_{\text{in}}}\mathbb{E}[M_{1,ij}^{2}]\cdot\mathbb{E}[M_{2,ij}^{2}]\cdot\mathbb{E}[x_{j}^{2}].$
To compute $\mathbb{E}[M_{1,ij}^{2}]$ and $\mathbb{E}[M_{2,ij}^{2}]$, we use the fact that the entries of $A_{i}$ and $B_{i}$ are i.i.d., yielding:
$\mathbb{E}[M_{l,ij}^{2}]=\sum_{k=1}^{r_{l}}\mathbb{E}[B_{l,ik}^{2}]\cdot\mathbb{E}[A_{l,kj}^{2}]=r_{l}\cdot\mathbb{E}[B_{l,11}^{2}]\cdot\mathbb{E}[A_{l,11}^{2}].$
Combining terms, we obtain:
$\mathbb{E}[y_{i}^{2}]$ $=s_{\text{ABBA}}^{2}\cdot d_{\text{in}}\cdot r_{1}r_{2}\cdot\mathbb{E}[B_{1,11}^{2}]\cdot\mathbb{E}[A_{1,11}^{2}]\cdot\mathbb{E}[B_{2,11}^{2}]\cdot\mathbb{E}[A_{2,11}^{2}]\cdot\mathbb{E}[x_{j}^{2}],$
$\mathbb{E}[(\nabla_{x})_{j}^{2}]$ $=s_{\text{ABBA}}^{2}\cdot d_{\text{out}}\cdot r_{1}r_{2}\cdot\mathbb{E}[B_{1,11}^{2}]\cdot\mathbb{E}[A_{1,11}^{2}]\cdot\mathbb{E}[B_{2,11}^{2}]\cdot\mathbb{E}[A_{2,11}^{2}]\cdot\mathbb{E}[(\nabla_{y})_{i}^{2}].$
Stability under initialization.
To simplify our analysis, we assume that instead of setting $B_{1}=\mathbf{0}$, its entries are drawn i.i.d. from a distribution $\mathcal{D}$ with variance $\sigma_{B_{1}}^{2}$. Given that $s_{\text{ABBA}}\propto\frac{1}{\sqrt{r_{1}r_{2}}}$, we find:
Forward $\text{2}^{\text{nd}}$ moment:
Assuming $\mathbb{E}[x_{j}^{2}]=\Theta_{r_{1},r_{2}}(1)$, we have:
$\mathbb{E}[y_{i}^{2}]=\mathcal{O}_{r_{1},r_{2}}(1)\cdot\sigma_{B_{1}}^{2}.$
Backward $\text{2}^{\text{nd}}$ moment:
Assuming $\mathbb{E}[(\nabla_{y})_{i}^{2}]=\Theta_{r_{1},r_{2}}(1)$, we similarly obtain:
$\mathbb{E}[(\nabla_{x})_{j}^{2}]=\mathcal{O}_{r_{1},r_{2}}(1)\cdot\sigma_{B_{1}}^{2}.$
Conclusion.
Following Definition 1, these results imply that ABBA adapters are rank-stabilized. Moreover, even in the original case where $B_{1}=\mathbf{0}$, we observe that as $\sigma_{B_{1}}\to 0$, both variances converge to $\Theta_{r_{1},r_{2}}(0)$, still satisfying the criteria for rank stability. Therefore, we conclude that ABBA adapters remain rank-stabilized under the given initialization and scaling. ∎
@@ Appendix E Selecting $\alpha$
In Theorem 2, we established the scaling relationship between ABBA and LoRA by comparing their respective factors, $\alpha_{\text{ABBA}}$ and $\alpha_{\text{LoRA}}$. To validate this empirically, we perform a sweep over different $\alpha_{\text{LoRA}}$ values and compute the corresponding ABBA scaling factors using $s_{\text{ABBA}}=\alpha_{\text{LoRA}}^{2}/\sqrt{r_{1}r_{2}}$, reporting results for Mistral-7B in Table 7. In line with our theoretical analysis and prior results in Table 4, ABBA performs best when $\alpha_{\text{LoRA}}$ lies in the typical range of 16–32.
Table 7: Performance comparison across different scaling factor values $s_{\text{ABBA}}$ for Mistral-7B.
$\mathbf{s_{\textbf{ABBA}}}$ ($\mathbf{\alpha_{\textbf{LoRA}}}$) Accuracy ($\uparrow$)
@@
$1$ ($4$) $62.17$ $17.10$
$4$ ($8$) $64.06$ $17.60$
$8$ ($8\sqrt{2}$) $64.43$ $18.14$
$16$ ($16$) $66.10$ $18.16$
$32$ ($16\sqrt{2}$) $66.26$ $18.08$
$64$ ($32$) $65.81$ $17.68$
$128$ ($32\sqrt{2}$) $65.20$ $16.82$
$256$ ($64$) $64.79$ $15.08$
Appendix F Varying Rank
@@ Appendix H Experimental Details
We implement all models using PyTorch [50] and HuggingFace Transformers [51]. All experiments run on a single NVIDIA A6000 GPU (48 GB). To reduce memory usage, we initialize base models in torch.bfloat16 precision. We train each configuration with the AdamW optimizer [52] and report the mean performance over three random seeds.
@@
While we adopt most settings from prior work [27], we perform a targeted learning rate sweep to optimize performance. For baseline comparisons, we replicate the experimental setups from the original PiSSA [26], rsLoRA [19], DoRA [11], LoRA-Pro [14], and HiRA [13] papers to ensure fair and consistent evaluation.
@@
Epochs $2$ $1$
Dropout $50.0$
Learning Rate $1\times 10^{-3}$ $1\times 10^{-3}$
@@ Appendix I Dataset Details
CommonSense170K is a unified benchmark that aggregates eight commonsense reasoning datasets into a single multi-task setting [27]. Each instance is a multiple-choice question, and models are prompted to select the correct answer without providing explanations. We adopt the prompt format introduced by the paper [27]. Below, we briefly describe the constituent datasets:
  • OBQA [28]: Open-book QA requiring retrieval and multi-hop reasoning over external knowledge.
  • ARC Challenge (ARC-c) [29]: Difficult grade-school science questions designed to test advanced reasoning beyond surface heuristics.
  • ARC Easy (ARC-e) [29]: Simpler science questions assessing core factual and conceptual understanding.
  • WinoGrande [30]: Pronoun resolution tasks requiring commonsense inference to resolve ambiguity.
  • HellaSwag [31]: Next-sentence prediction under a constrained completion setting, testing grounded understanding of everyday scenarios.
  • PIQA [32]: Physical reasoning tasks where models select the most sensible solution to a practical problem.
  • SIQA [33]: Social reasoning benchmark involving questions about intent, social dynamics, and consequences of human actions.
  • BoolQ [34]: Binary (yes/no) questions drawn from natural queries, requiring contextual understanding of short passages.
MetaMathQA [35] reformulates existing mathematical problems into alternative phrasings that preserve their original semantics, offering diverse surface forms without introducing new information. We evaluate models fine-tuned on this dataset using two benchmarks: GSM8K [36], which targets step-by-step reasoning in elementary arithmetic word problems, and MATH [37], which features high-difficulty problems drawn from math competitions. We evaluate solely based on the correctness of the final numeric answer.

### v3

@@ D.2 Proof of Theorem 2: Rank-Stability of ABBA
For any input $x\in\mathbb{R}^{d_{\text{in}}}$, the ABBA adapter produces
$y_{i}=s_{\text{ABBA}}\sum_{j=1}^{d_{\text{in}}}M_{1,ij}M_{2,ij}\cdot x_{j},\quad\forall i\in\{1,\dots,d_{\text{out}}\},$ (14)
where $M_{1}=B_{1}A_{1}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ and $M_{2}=B_{2}A_{2}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$. Thus
@@
The gradient with respect to the input is
$(\nabla_{x}L)_{j}$ $=\sum_{i=1}^{d_{\text{out}}}\frac{\partial L}{\partial y_{i}}\frac{\partial y_{i}}{\partial x_{j}}=s_{\text{ABBA}}\sum_{i=1}^{d_{\text{out}}}M_{1,ij}M_{2,ij}\,(\nabla_{y}L)_{i},$
where $(\nabla_{y}L)_{i}=\frac{\partial L}{\partial y_{i}}$. Hence
$\mathbb{E}[(\nabla_{x}L)_{j}^{2}]=s_{\text{ABBA}}^{2}\cdot\mathbb{E}\!\left[\left(\sum_{i=1}^{d_{\text{out}}}M_{1,ij}M_{2,ij}(\nabla_{y}L)_{i}\right)^{2}\right].$
Second-moment factorization.
Since $M_{1}$ and $M_{2}$ are independent with mean-zero entries,
$\mathbb{E}[M_{1,ij}^{2}]=r_{1}\,\mathbb{E}[B_{1,11}^{2}]\mathbb{E}[A_{1,11}^{2}],\quad\mathbb{E}[M_{2,ij}^{2}]=r_{2}\,\mathbb{E}[B_{2,11}^{2}]\mathbb{E}[A_{2,11}^{2}].$
Therefore
$\mathbb{E}[M_{1,ij}^{2}M_{2,ij}^{2}]=\mathbb{E}[M_{1,ij}^{2}]\,\mathbb{E}[M_{2,ij}^{2}]=r_{1}r_{2}\cdot\mathbb{E}[B_{1,11}^{2}]\mathbb{E}[A_{1,11}^{2}]\mathbb{E}[B_{2,11}^{2}]\mathbb{E}[A_{2,11}^{2}].$
Forward and backward moments.
Substituting,
$\mathbb{E}[y_{i}^{2}]$ $=s_{\text{ABBA}}^{2}\cdot d_{\text{in}}\cdot r_{1}r_{2}\cdot C\cdot\mathbb{E}[x_{j}^{2}],$
$\mathbb{E}[(\nabla_{x}L)_{j}^{2}]$ $=s_{\text{ABBA}}^{2}\cdot d_{\text{out}}\cdot r_{1}r_{2}\cdot C\cdot\mathbb{E}[(\nabla_{y}L)_{i}^{2}],$
where $C=\mathbb{E}[B_{1,11}^{2}]\mathbb{E}[A_{1,11}^{2}]\mathbb{E}[B_{2,11}^{2}]\mathbb{E}[A_{2,11}^{2}]$.
Rank-stability.
By Definition 1, forward stability requires $\mathbb{E}[y_{i}^{2}]=\Theta_{r_{1},r_{2}}(1)$ whenever $\mathbb{E}[x_{j}^{2}]=\Theta(1)$, and backward stability requires $\mathbb{E}[(\nabla_{x}L)_{j}^{2}]=\Theta_{r_{1},r_{2}}(1)$ whenever $\mathbb{E}[(\nabla_{y}L)_{i}^{2}]=\Theta(1)$. Both hold exactly when
$s_{\text{ABBA}}^{2}r_{1}r_{2}=\Theta(1),$
that is,
$s_{\text{ABBA}}\in\Theta\!\left(\tfrac{1}{\sqrt{r_{1}r_{2}}}\right).$
Following Definition 1, these results imply that ABBA adapters are rank-stabilized.
@@ Appendix E Selecting $\alpha$
We perform a sweep over different $\alpha$ values, reporting results for Mistral-7B in Table 7. In line with prior results in Table 4, ABBA performs best when $\alpha_{\text{LoRA}}$ lies in the typical range of 16–32.
Table 7: Performance comparison across different $\alpha$ values for Mistral-7B.
$\alpha$ Accuracy ($\uparrow$)
@@
44 $62.17$ $17.10$
88 $64.06$ $17.60$
$12$ $64.43$ $18.14$
$16$ $66.10$ $18.16$
$24$ $66.26$ $18.08$
$32$ $65.81$ $17.68$
$48$ $65.20$ $16.82$
$64$ $64.79$ $15.08$
Appendix F Effect of Varying Rank
@@ Appendix H Experimental Details
We implement all models using PyTorch (49) and HuggingFace Transformers (50). All experiments run on a single NVIDIA A6000 GPU (48 GB). To reduce memory usage, we initialize base models in torch.bfloat16 precision. We train each configuration with the AdamW optimizer (51) and report the mean performance over three random seeds.
@@
While we adopt most settings from prior work (27), we perform a targeted learning rate sweep to optimize performance. For baseline comparisons, we replicate the experimental setups from the original PiSSA (26), rsLoRA (21), DoRA (11), LoRA-Pro (14), and HiRA (13) papers to ensure fair and consistent evaluation.
@@
Epochs 22 11
Dropout $0.05$ 0
Learning Rate $\times 10^{-3} 1×1$$\times 10^{-3}$
@@ Appendix I Dataset Details
CommonSense170K is a unified benchmark that aggregates eight commonsense reasoning datasets into a single multi-task setting (27). Each instance is a multiple-choice question, and models are prompted to select the correct answer without providing explanations. We adopt the prompt format introduced by the paper (27). Below, we briefly describe the constituent datasets:
  • OBQA (28): Open-book QA requiring retrieval and multi-hop reasoning over external knowledge.
  • ARC Challenge (ARC-c) (29): Difficult grade-school science questions designed to test advanced reasoning beyond surface heuristics.
  • ARC Easy (ARC-e) (29): Simpler science questions assessing core factual and conceptual understanding.
  • WinoGrande (30): Pronoun resolution tasks requiring commonsense inference to resolve ambiguity.
  • HellaSwag (31): Next-sentence prediction under a constrained completion setting, testing grounded understanding of everyday scenarios.
  • PIQA (32): Physical reasoning tasks where models select the most sensible solution to a practical problem.
  • SIQA (33): Social reasoning benchmark involving questions about intent, social dynamics, and consequences of human actions.
  • BoolQ (34): Binary (yes/no) questions drawn from natural queries, requiring contextual understanding of short passages.
MetaMathQA (35) reformulates existing mathematical problems into alternative phrasings that preserve their original semantics, offering diverse surface forms without introducing new information. We evaluate models fine-tuned on this dataset using two benchmarks: GSM8K (36), which targets step-by-step reasoning in elementary arithmetic word problems, and MATH (37), which features high-difficulty problems drawn from math competitions. We evaluate solely based on the correctness of the final numeric answer.
Appendix J Use of Large Language Models
LLMs are used exclusively for minor writing assistance, for example, to enhance grammar and improve sentence clarity.